# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

### v1

@@ Abstract
We develop and analyze a stochastic genetic interacting particle method (SGIP) for reaction-diffusion-advection (RDA) equations. The method employs operator splitting to approximate the advection-diffusion and reaction processes, treating the former by particle drift-diffusion and the latter by exact or implicit integration of reaction dynamics over bins where particle density is estimated by a histogram. A key innovation is the incorporation of adaptive resampling to close the loop of particle and density field description of solutions, mimicking mutation in genetics. Resampling is also critical for maintaining long-term stability by redistributing particles with the evolving density field. We provide a comprehensive error analysis, and establish rigorous convergence bounds under appropriate regularity assumptions. Numerical experiments from one to three space dimensions demonstrate the method’s effectiveness across various reaction types (FKPP, cubic, Arrhenius) and flow configurations (shear, cellular, cat’s eye, ABC flows), showing excellent agreement with finite difference method (FDM) while offering computational advantages for complex flow geometries and higher-dimensional problems.
Keywords: Stochastic particle method, resampling, reaction-diffusion-advection equations, operator splitting, convergence analysis, 3D simulations.
@@ 1 Introduction
Reaction-Diffusion-Advection (RDA) equations play a fundamental role in modeling numerous physical, biological, and chemical processes. The general form of RDA equations is given by:
$\frac{\partial u}{\partial t}+\nabla\cdot(\mathbf{v}u)=D\Delta u+r(u)$ (1)
@@
where $u(\mathbf{x},t)$ represents the concentration or density field, $\mathbf{v}(\mathbf{x},t)$ is ambient fluid flow field, $D$ is molecular diffusion coefficient, and $r(u)$ is reaction nonlinearity. The versatility of RDA equations is reflected in their wide-ranging applications [6, 24, 35]. In ecology, the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation models propagating population and gene waves [11], while developmental biology employs Turing’s reaction-diffusion model to explain morphological patterning [33]. The framework also extends to materials science for simulating phase separation via the Cahn-Hilliard equation [4], and to finance, where it underpins the Black-Scholes model for option pricing [3], collectively underscoring its broad utility [18].
@@
The development of efficient numerical methods for these equations remains an active research area, with challenges from nonlinearities, multiscale phenomena, and high-dimensionality. The Finite Difference Method (FDM) is well-suited for simple geometries, with established stability and convergence properties [21, 14]. For complex domains, the Finite Element Method (FEM) offers flexibility, grounded in variational principles [20, 17, 28]. Operator splitting techniques (e.g., Lie–Trotter and Strang splitting) are often used to mitigate stiffness from diffusion or reaction terms [32, 31]. Despite their successes, the mesh-based methods struggle with the curse of dimensionality and problems involving large deformations or moving boundaries where mesh maintenance becomes prohibitively expensive.
@@
Particle (Lagrangian) techniques present a mesh-free alternative to handle large deformation (e.g. sharp internal layers) and high-dimensional problems. In the 1980’s, [29] [29, 30] introduced a random walk method on the spatial derivative of solutions. The reaction is treated by a reproduction and destruction process to evolve steep wavefronts with an application to Hodgkin–Huxley equations. Representing spatial derivative by an empirical measure of particles, Puckett [26] showed its convergence for the FKPP equation. These early works are limited to one space (1D) dimension and monotone initial data however. For recent efforts on particle methods of RDA type equations, we refer to [22, 19, 37, 23] and references therein. In general, a particle description alone can be either computationally expensive or insufficient to represent solutions of physical and biological systems. In the absence of low order reactions, the current authors [34, 15, 16] recently introduced efficient stochastic particle-field methods to compute parabolic-parabolic chemotaxis and haptotaxis advection-diffusion systems up to 3D, in lieu of a history dependent pure particle description involving heat kernels.
In the spirit of particle-field representation, the stochastic genetic interacting particle (SGIP) method developed in this work provides an efficient and robust framework for RDA equations in several space dimensions. It is a natural extension of the SGIP [22, 37] built from the Feynman-Kac semigroup for linear RDA equations. The algorithm consists of a 3-step operator-splitting to decompose the RDA process: drift-diffusion, density update and particle resampling. The advection-diffusion is handled by drift-diffusion (Ito equation) of particles. The reaction process is approximated by going between particles and density: the density field is a histogram counting particle positions over bins, and is evolved in time via reaction ODEs by either an exact (e.g. for FKPP) or a stable implicit integration (Algorithm 2). Then particles are redrawn from the updated density by an adaptive resampling step (Algorithm 3) so that they are ready for advection-diffusion again to close the loop. The joint particle-density evolution is a way to overcome the difficulty of seeking a self-contained particle description. In case of SGIP for linear RDA [22, 37], the particle dynamics is in a close loop via drift-diffusion and resampling steps without relying on density field, due to reaction rate being a known potential function. The resampling step here mimics mutation in genetics in that particles are more likely to survive in a bin to the next generation if the bin has more particles (is more healthy) than its neighbors. Resampling is also critical for long-term stability and computational efficiency by maintaining a well-balanced representation of the evolving solution, i.e. keeping variance in check over time. Our SGIP framework here is supported by a rigorous convergence analysis under appropriate regularity assumptions, which establishes consistent accuracy controlled by time step, spatial resolution, and the number of particles.
Our numerical experiments comprehensively demonstrate SGIP’s capabilities across various dimensions and complexities. In 1D scenarios, SGIP shows excellent agreement with FDM for different reaction types (FKPP, cubic, and Arrhenius) [10, 1], accurately capturing propagation speeds and front shapes. The method effectively resolves a spectrum of complex flow configurations, from fundamental 2D shear, cellular and cat’s eye flows to challenging 3D Arnold-Beltrami-Childress (ABC) flows [27, 7, 25, 9, 36], even when coupled with Lipschitz-continuous reaction terms. In these scenarios, particularly in 3D, SGIP maintains high accuracy while offering significant computational advantages over traditional mesh-based methods like the FDM. As the diffusion coefficient decreases and advection dominates, the SGIP framework remains stable and efficient, in stark contrast to FDM, which suffers from prohibitive computational costs and severe numerical instability under such conditions.
@@
The SGIP method provides a simple-to-program, scalable, low-cost and stable framework for high-dimensional RDA problems. Its mesh-free formulation bypasses the geometric constraints of grid-based methods, allowing it to handle complex flows and large gradients of solutions with ease. This capability is augmented by favorable computational scaling in high dimensions and long-term stability, thanks to adaptive resampling. These combined strengths allow SGIP to open new avenues for simulating complex multiscale phenomena in physical and biological systems, particularly in scenarios involving intricate flow geometries and requiring long-time integration.
@@
The rest of the paper is organized as follows. In Section 2, we present the detailed formulation of the SGIP algorithm. This section is divided into three subsections dedicated to the stochastic advection-diffusion step, the reaction step, and the adaptive resampling strategy, respectively. Section 3 is devoted to the theoretical analysis, where a comprehensive error decomposition is provided and convergence bounds are rigorously established. In Section 4, extensive numerical experiments are conducted to validate the method’s performance, ranging from 1D benchmark tests to 2D and 3D simulations involving complex flows. Finally, we conclude with a summary of our findings and a discussion of future research directions in Section 5.
@@ 2 SGIP Algorithm
In this section, we will present the SGIP algorithm for solving the reaction-diffusion-advection (RDA) equations. The mathematical foundation of our method is the operator-splitting technique. We consider the RDA equation as an abstract evolution equation:
$\frac{\partial u}{\partial t}=(\mathcal{A}+\mathcal{R})u,\quad u(0)=u_{0}$ (2)
@@
$u_{n+1}\approx\mathcal{S}_{\mathcal{R}}^{\Delta t}\circ\mathcal{S}_{\mathcal{A}}^{\Delta t}(u_{n})$ (3)
Here, $u_{n}$ denotes the numerical approximation of $u(\mathbf{x},t_{n})$ at time $t_{n}$, and $u_{n+1}$ denotes the approximation at the next time level $t_{n+1}=t_{n}+\Delta t$. This splitting enables us to employ specialized numerical techniques for each physical process, ensuring its long-term stability and efficiency.
For the advection-diffusion part $\mathcal{S}_{\mathcal{A}}^{\Delta t}$, we employ the stochastic particle method, which turns out to be stable in various high-dimensional advection-dominant problems. For the reaction $\mathcal{S}_{\mathcal{A}}^{\Delta t}$ counterpart, we apply a density-based approach, which recovers the density $u$ from the particle formulation and approximates the reaction $\mathcal{R}$ by resampling. Details of each part of the algorithm are listed in the following subsections.
@@ 2.1 Initialization
$u(\mathbf{x},t)\approx\frac{M_{0}}{N}\sum_{i=1}^{N}\delta(\mathbf{x}-\mathbf{X}^{i}(t))$ (4)
@@
$\mathbf{X}_{0}^{i}\sim\rho_{0}(\mathbf{x})=\frac{u_{0}(\mathbf{x})}{M_{0}},\quad i=1,\dots,N$ (5)
Particle masses are initialized uniformly according to:
$m_{0}^{i}=\frac{M_{0}}{N},\quad i=1,\dots,N.$ (6)
In the temporal direction, we assume that the computational time range $[0,T]$ is partitioned into time steps by $\{t_{n}\}_{n=0}^{n_{T}}$, where $t_{0}=0$ and $t_{n_{T}}=T$. We consider the uniform time step, i.e., there exists $\Delta t=t_{n+1}-t_{n}$ for $n=0,\dots,n_{T}-1$. For ease of presenting our algorithm, we will use a slight abuse of notation. We will represent the density $u$ at time $t_{n}$ as $u_{n}$, and the particle positions at time $t_{n}$ as $\{\mathbf{X}_{n}^{i}\}_{i=1}^{N}$. Similarly, particle masses at time $t_{n}$ are denoted as $m_{n}^{i}$.
@@ 2.2 Advection-Diffusion Step
The advection-diffusion sub-step $\mathcal{S}_{\mathcal{A}}^{\Delta t}$ advances the solution from $t_{n}$ to an intermediate time. Each particle follows the stochastic differential equation:
$d\mathbf{X}^{i}=\mathbf{v}(\mathbf{X}^{i},t)dt+\sqrt{2D}d\mathbf{W}^{i}(t)$ (7)
@@
$\mathbf{X}_{*}^{i}=\mathbf{X}_{n}^{i}+\mathbf{v}(\mathbf{X}_{n}^{i},t_{n})\Delta t+\sqrt{2D\Delta t}\boldsymbol{\xi}_{n}^{i}$ (8)
where $\boldsymbol{\xi}_{n}^{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ are independent standard normal random variables, and $\mathbf{X}_{*}^{i}$ denotes the intermediate particle positions.
@@ 2.3 Reaction Step
The reaction sub-step $\mathcal{S}_{\mathcal{R}}^{\Delta t}$ handles the nonlinear reaction term from the intermediate state to time $t_{n+1}$. This step consists of two main components: density estimation on a Eulerian grid, followed by reaction integration.
@@ 2.3.1 Density Estimation
To evaluate the reaction terms, we first reconstruct the density field from the particle distribution. The domain $\Omega=[-L,L]^{d}$ is discretized into $M^{d}$ uniform bins with centers $\{\mathbf{x}_{j}\}_{j=1}^{M^{d}}$ and bin size $\Delta x=2L/M$. The density in each bin is estimated by counting particles and summing their masses:
$\hat{u}_{*}^{j}=\frac{1}{(\Delta x)^{d}}\sum_{i:\mathbf{X}_{*}^{i}\in\text{bin}_{j}}m_{n}^{i}$ (9)
@@ 2.3.2 Reaction Integration
For certain reaction terms, exact integration is possible. A prominent example is the FKPP reaction:
r(u)=u(1u)r(u)=u(1-u) (10)
which admits the exact solution:
$\hat{u}_{n+1}^{j}=\frac{\hat{u}_{*}^{j}e^{\Delta t}}{1+\hat{u}_{*}^{j}(e^{\Delta t}-1)}$ (11)
Other examples include linear reactions $r(u)=\lambda u$, which yield exponential growth/decay $\hat{u}_{n+1}^{j}=\hat{u}_{*}^{j}e^{\lambda\Delta t}$.
@@
For reaction terms without closed-form solutions, such as:
@@
  • Polynomial reactions: $r(u)=\sum_{k=0}^{K}a_{k}u^{k}$,
we employ implicit time integration methods to maintain numerical stability. The backward Euler scheme reads:
$\frac{\hat{u}_{n+1}^{j}-\hat{u}_{*}^{j}}{\Delta t}=r(\hat{u}_{n+1}^{j})$ (12)
which is solved numerically using Newton’s method. For stiff problems, higher-order implicit methods such as Crank-Nicolson [8] or BDF2 [12] can be employed.
The reaction integration is performed independently in each spatial bin, making this step highly parallelizable. After reaction integration, we obtain the updated density field $\{\hat{u}_{n+1}^{j}\}$ at time $t_{n+1}$.

### v2

@@ Abstract
We develop and analyze a stochastic genetic interacting particle method (SGIP) for reaction-diffusion-advection (RDA) equations. The SGIP method employs operator splitting to approximate the advection-diffusion and reaction processes, treating the former using particle drift-diffusion and the latter via exact or implicit integration of reaction dynamics over bins, where particle density is estimated using a histogram. A key innovation is the incorporation of adaptive resampling to close the loop of particle and density field description of solutions, mimicking the selection mechanism in genetics. Resampling is also crucial for maintaining long-term stability by redistributing particles in accordance with the evolving density field. We provide a comprehensive error analysis and establish convergence bounds under appropriate regularity assumptions. Numerical experiments in one to three space dimensions demonstrate the method’s effectiveness across various reaction types (Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP), cubic, Arrhenius) and flow configurations (shear, cellular, cat’s eye, Arnold-Beltrami-Childress (ABC) flows), showing excellent agreement with the finite difference method (FDM) while offering computational advantages for complex flow geometries and higher-dimensional problems.
keywords:
Stochastic particle method, resampling methods, reaction-diffusion-advection (RDA) equations, operator splitting, convergence analysis, 3D simulations.
{MSCcodes}
35K57, 65C35, 65M12, 65M75, 65J08.
@@ 1 Introduction
Reaction-Diffusion-Advection (RDA) equations play a fundamental role in modeling numerous physical, biological, and chemical processes. The general form of RDA equations is given by
(1) $\frac{\partial u}{\partial t}+\nabla\cdot(\mathbf{v}u)=D\Delta u+r(u)$
@@
where $u(\mathbf{x},t)$ represents the concentration or density field, $\mathbf{v}(\mathbf{x},t)$ is the ambient fluid flow field, $D$ is the molecular diffusion coefficient, and $r(u)$ is the reaction nonlinearity. The versatility of RDA equations is reflected in their wide-ranging applications [chorin1973numerical, peskin1977numerical, williams2018combustion]. In ecology, the Fisher-Kolmogorov-Petrovsky-Piskunov (FKPP) equation models the propagation of population and gene waves [fisher1937wave], whereas developmental biology employs Turing’s reaction-diffusion model to explain morphological patterning [turing1952chemical]. The framework also extends to materials science for simulating phase separation via the Cahn-Hilliard equation [cahn1958free], and to finance, where it underpins the Black-Scholes model for option pricing [black1973pricing], collectively underscoring its broad utility [hundsdorfer2013numerical].
@@
The development of efficient numerical methods for these equations remains an active research area, with challenges from nonlinearities, multiscale phenomena, and high-dimensionality. The finite difference method (FDM) is well-suited for simple geometries, with established stability and convergence properties [leveque2007finite, hoff1978stability]. For complex domains, the Finite Element Method (FEM) offers flexibility, grounded in variational principles [komala2018finite, hughes2012finite, shen2013finite]. Operator splitting techniques (e.g., Lie-Trotter and Strang splitting) are often used to mitigate stiffness from diffusion or reaction terms [trotter1959product, strang1968construction]. Despite their successes, the mesh-based methods struggle with the curse of dimensionality and problems involving large deformations or moving boundaries, forcing the computationally expensive and error-prone process of repeated mesh regeneration to avoid severe element distortion.
@@
Particle (Lagrangian) techniques present a mesh-free alternative to handle large deformation (e.g., sharp internal layers) and high-dimensional problems. In the 1980s, Sherman and Peskin [sherman1986monte, sherman1988solving] introduced a random walk method on the spatial derivative of solutions. The reaction is treated through a reproduction and destruction process to evolve steep wavefronts, with applications to Hodgkin-Huxley equations. Puckett [puckett1989convergence] proved the convergence of a random gradient method for the FKPP equation, where the solution gradient is represented by an empirical measure of particles with weights corresponding to jumps in a step-function approximation. These early works are limited to the one-dimensional (1D) case and to monotone initial data, however. For recent efforts on particle methods of RDA-type equations, we refer to [lyu2022convergent, ParWasser_2025, zhang2025convergent, lyu2025stochastic] and references therein. In the absence of low-order reactions, the current authors [wang2025novel, hu2024stochastic, hu2025convergence] recently introduced efficient stochastic particle-field methods to compute parabolic-parabolic chemotaxis and haptotaxis advection-diffusion systems up to 3D, instead of a history-dependent pure particle description involving heat kernels.
In the spirit of particle-field representation, the stochastic genetic interacting particle (SGIP) method developed in this work provides an efficient and robust framework for RDA equations in several space dimensions. The method naturally extends the SGIP framework [lyu2022convergent, zhang2025convergent], originally constructed from the Feynman-Kac semigroup for linear RDA equations, to handle the nonlinear and multi-dimensional settings. The algorithm consists of a 3-step operator-splitting to decompose the RDA process: drift-diffusion, density update, and particle resampling.
The drift-diffusion process of the particles handles the advection-diffusion part of the PDE system through its probabilistic representation (Algorithm 1). The reaction process is approximated by going between particles and density: we reconstruct the density field by the particle locations, and evolve in time via reaction ODEs by either an exact (e.g., for FKPP) or a stable implicit integration (Algorithm 2). Then, particles are re-drawn from the updated density using an adaptive resampling step (Algorithm 3) so that they are ready for advection-diffusion again, closing the loop. The joint particle-density evolution is a way to overcome the difficulty of seeking a self-contained particle description. As a comparison, SGIP for linear RDA [lyu2022convergent, zhang2025convergent], the particle dynamics are in a closed loop via drift-diffusion and resampling steps without relying on the density field reconstruction.
@@
The SGIP method provides a simple-to-program, scalable, low-cost, and stable framework for high-dimensional RDA problems. Its mesh-free formulation bypasses the geometric constraints of grid-based methods, allowing it to handle complex flows and large gradients of solutions with ease. This capability is augmented by favorable computational scaling in high dimensions and long-term stability, thanks to adaptive resampling. Furthermore, the SGIP framework here is supported by a complete convergence analysis under appropriate regularity assumptions, which establishes consistent accuracy controlled by time step, spatial resolution, and the number of particles.
Our numerical experiments comprehensively demonstrate SGIP’s capabilities across various dimensions and complexities, revealing its superior computational efficiency—particularly in demanding 3D settings where traditional mesh-based methods become prohibitively expensive. In 1D scenarios, SGIP shows excellent agreement with FDM for different reaction types (FKPP, cubic, and Arrhenius) [fife2013mathematical, arrhenius1889reaktionsgeschwindigkeit], accurately capturing propagation speeds and front shapes. The method effectively resolves a spectrum of complex flow configurations, from fundamental 2D shear, cellular, and cat’s eye flows to challenging 3D Arnold-Beltrami-Childress (ABC) flows [ranz1979applications, constantin2000bulk, pierrehumbert1991large, dombre1986chaotic, xin2000front], even when coupled with Lipschitz-continuous reaction terms. In these scenarios, particularly in 3D, SGIP maintains high accuracy while offering significant computational advantages over traditional mesh-based methods, such as the FDM. As the diffusion coefficient decreases and advection dominates, the SGIP framework remains stable and efficient, in stark contrast to FDM, which suffers from prohibitive computational costs and severe numerical instability under such conditions.
@@
The rest of the paper is organized as follows. In Section 2, we present the detailed formulation of the SGIP algorithm. This section is divided into three subsections dedicated to the stochastic advection-diffusion step, the reaction step, and the adaptive resampling strategy, respectively. Section 3 is devoted to the theoretical analysis, where a comprehensive error decomposition is provided, and convergence bounds are established. In Section 4, extensive numerical experiments are conducted to validate the method’s performance, ranging from 1D benchmark tests to 2D and 3D simulations involving complex flows. Finally, we conclude with a summary of our findings and a discussion of future research directions in Section 5.
@@ 2 SGIP Algorithm
In this section, we will present the SGIP algorithm for solving the reaction-diffusion-advection (RDA) equations. The mathematical foundation of our method is the operator-splitting technique. In particular, we consider the RDA equation as an abstract evolution equation:
(2) $\frac{\partial u}{\partial t}=(\mathcal{A}+\mathcal{R})u,\quad u(0)=u_{0},$
@@
(3) $u_{n+1}\approx\mathcal{S}_{\mathcal{R}}(\Delta t)\circ\mathcal{S}_{\mathcal{A}}(\Delta t)(u_{n}).$
Here, $u_{n}$ denotes the numerical approximation of $u(\mathbf{x},t_{n})$ at time $t_{n}$, and $u_{n+1}$ denotes the approximation at the next time level $t_{n+1}=t_{n}+\Delta t$. The operators $\mathcal{S}_{\mathcal{A}}(\Delta t)$ and $\mathcal{S}_{\mathcal{R}}(\Delta t)$ are numerical approximations over a single time step $\Delta t$ to the exact solutions of the subproblems $\partial_{t}u=\mathcal{A}u$ and $\partial_{t}u=\mathcal{R}u$, respectively. This splitting enables us to employ specialized numerical techniques for each physical process, ensuring its long-term stability and efficiency.
For the advection-diffusion part $\mathcal{S}_{\mathcal{A}}(\Delta t)$, we employ the stochastic particle method, which turns out to be stable in various high-dimensional advection-dominant problems. For the reaction counterpart $\mathcal{S}_{\mathcal{R}}({\Delta t})$, we apply a density-based approach, which recovers the density $u$ from the particle formulation and approximates the reaction term by resampling. Details of each part of the algorithm are listed in the following subsections.
@@ 2.1 Initialization
(4) $u(\mathbf{x},t)\approx\frac{M_{0}}{N}\sum_{i=1}^{N}\delta(\mathbf{x}-\mathbf{X}^{i}(t)),$
@@
(5) $\mathbf{X}_{0}^{i}\sim\rho_{0}(\mathbf{x})=\frac{u_{0}(\mathbf{x})}{M_{0}},\quad i=1,\dots,N.$
Particle masses are initialized uniformly according to
(6) $m_{0}^{i}=\frac{M_{0}}{N},\quad i=1,\dots,N.$
In the temporal direction, we assume that the computational time range $[0,T]$ is partitioned into time steps defined by $\{t_{n}\}_{n=0}^{n_{T}}$, where $t_{0}=0$ and $t_{n_{T}}=T$. We consider the uniform time step, i.e., there exists $\Delta t=t_{n+1}-t_{n}$ for $n=0,\dots,n_{T}-1$. For ease of presenting our algorithm, we will use a slight abuse of notation. We will represent the density $u$ at time $t_{n}$ as $u_{n}$, and the particle positions at time $t_{n}$ as $\{\mathbf{X}_{n}^{i}\}_{i=1}^{N}$. Similarly, particle masses at time $t_{n}$ are denoted as $m_{n}^{i}$.
@@ 2.2 Advection-Diffusion Step
The advection-diffusion sub-step $\mathcal{S}_{\mathcal{A}}({\Delta t})$ advances the solution from $t_{n}$ to an intermediate time. Each particle follows the stochastic differential equation:
(7) $d\mathbf{X}^{i}=\mathbf{v}(\mathbf{X}^{i},t)dt+\sqrt{2D}d\mathbf{W}^{i}(t),$
@@
(8) $\mathbf{X}_{*}^{i}=\mathbf{X}_{n}^{i}+\mathbf{v}(\mathbf{X}_{n}^{i},t_{n})\Delta t+\sqrt{2D\Delta t}\boldsymbol{\xi}_{n}^{i},$
where $\boldsymbol{\xi}_{n}^{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ are independent standard normal random variables, and $\mathbf{X}_{*}^{i}$ denotes the intermediate particle positions. The complete procedure for this step is summarized in Algorithm 1.
@@ 2.3 Reaction Step
The reaction sub-step $\mathcal{S}_{\mathcal{R}}(\Delta t)$ handles the nonlinear reaction term from the intermediate state to time $t_{n+1}$. This step consists of two main components: density estimation on an Eulerian grid, followed by reaction integration.
@@ 2.3.1 Density Estimation
To evaluate the reaction terms, we first reconstruct the density field from the particle distribution onto a uniform Cartesian grid covering the domain $\Omega=[-L,L]^{d}$. Let $K$ be the number of grid intervals in each spatial direction, giving a cell size of $\Delta x=2L/K$. We define the index set of all grid cells as
(9) $\mathcal{G}:=\{1,\dots,K\}^{d}.$
The cell centers, denoted by $\{\mathbf{x}_{j}\}_{j\in\mathcal{G}}$, serve as the reconstruction points. The density $\{\hat{u}_{*}^{j}\}_{j\in\mathcal{G}}$ in each bin is estimated by counting particles and summing their masses:
(10) $\hat{u}_{*}^{j}=\frac{1}{(\Delta x)^{d}}\sum_{i:\mathbf{X}_{*}^{i}\in\text{bin}_{j}}m_{n}^{i}.$
@@ 2.3.2 Reaction Integration
For certain reaction terms, exact integration is possible. A prominent example is the FKPP reaction: r(u)=u(1u)r(u)=u(1-u), which admits the exact solution: $\hat{u}_{n+1}^{j}=\frac{\hat{u}_{*}^{j}e^{\Delta t}}{1+\hat{u}_{*}^{j}(e^{\Delta t}-1)}$. Other examples include linear reactions $r(u)=\lambda u$, which yield exponential growth or decay $\hat{u}_{n+1}^{j}=\hat{u}_{*}^{j}e^{\lambda\Delta t}$.
@@
For reaction terms without closed-form solutions:
@@
an implicit time discretization is adopted to maintain numerical stability for stiff reaction dynamics. The backward Euler scheme reads:
(11) $\frac{\hat{u}_{n+1}^{j}-\hat{u}_{*}^{j}}{\Delta t}=r(\hat{u}_{n+1}^{j}),$
which is solved numerically using Newton’s method. For stiff problems, higher-order implicit methods such as Crank-Nicolson [crank1947practical] or BDF2 [gear1971numerical] can be employed.
The reaction integration is performed independently in each spatial bin, making this step highly parallelizable. After reaction integration, we obtain the updated density field $\{\hat{u}_{n+1}^{j}\}_{j\in\mathcal{G}}$ at time $t_{n+1}$. The full workflow for the reaction step is detailed in Algorithm 2.
2.4 Resampling Step
Following the reaction integration that produces the updated density field $\{\hat{u}_{n+1}^{j}\}_{j\in\mathcal{G}}$, we employ a resampling strategy to redistribute particles according to this new density distribution. This step ensures particle efficiency and adaptivity while maintaining mass conservation.